OpenStudy (anonymous):

Can anyone please explain what properties of real numbers are use in this problem. thanks! 4x+3=0 4x+3+(-3) = 0 + (-3) 4x+(3-3) = -3 4x+0 = -3 4x * 1/4 = -3 * 1/4 x * 4 * 1/4 = 3/4 x* (4 * 1/4) = -3/4 x * 1 = -3/4 x = -3/4

OpenStudy (goformit100):

4x+3=0 : Multiplicative property 4x+3+(-3) = 0 + (-3) : Commutitivity and distributivity 4x+(3-3) = -3 : Distributive 4x+0 = -3 : additive Identity 4x * 1/4 = -3 * 1/4 : Reciprocal( Multiplicative Identity) x * 4 * 1/4 = 3/4 : Multiplication x* (4 * 1/4) = -3/4 : Distributivity x * 1 = -3/4 : Multiplicative Identity x = -3/4 : Value Denoting

OpenStudy (goformit100):

@shubhamsrg Sir can you please check if it is wrong or not.....?

OpenStudy (shubhamsrg):

am sorry am not too good with the names of properties. but perhaps you are right

OpenStudy (goformit100):

Thankyou sir

OpenStudy (skullpatrol):

4x+3=0_Given 4x+3+(-3) = 0 + (-3)_Addition Property of Equality 4x+(3-3) = -3_Definition of Subtraction, Identity Property of Addition 4x+0 = -3_Substitution Principle 4x * 1/4 = -3 * 1/4_Identity Property of Addition, Multiplication Property of Equality x * 4 * 1/4 = -3/4_Commutative Property of Multiplication, Definition of Division x* (4 * 1/4) = -3/4_Associative Property of Multiplication x * 1 = -3/4_Property of Reciprocals x = -3/4_Identity Property of Multiplication

OpenStudy (anonymous):

thanks guys! but which one is the right answer? im not good at naming real number properties too. hehe

OpenStudy (skullpatrol):

I am sure my answer is correct :)